OpenStudy (anonymous):

What is the slope of the line 3x + 4y = -11?

geerky42 (geerky42):

Can you get it to slope-intercept form: y = mx+b?

OpenStudy (anonymous):

y=-3/4x+-11/4

OpenStudy (anonymous):

m=-3/4 **